caydenisdead:

Solve 3(2x - 1) = 4x + 5

IMURDADDY4EVER:

OK, we going step by step. What is the first thing you need to do?

caydenisdead:

idk

caydenisdead:

evaluate

IMURDADDY4EVER:

do you know PEMDAS

caydenisdead:

no-

caydenisdead:

like i said SLOW

XxDaddy4ever:

Start wit the ()

sxdsouls:

Step 1: Simplify both sides of the equation. 3(2x−1)=4x+5 (3)(2x)+(3)(−1)=4x+5(Distribute) 6x+−3=4x+5 6x−3=4x+5 Step 2: Subtract 4x from both sides. 6x−3−4x=4x+5−4x 2x−3=5 Step 3: Add 3 to both sides. 2x−3+3=5+3 2x=8 Step 4: Divide both sides by 2. 2x 2 = 8 2 x=4 x=4 is the answer. also how many accounts do you have Cayden bruh..-

XxDaddy4ever:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sxdsouls Step 1: Simplify both sides of the equation. 3(2x−1)=4x+5 (3)(2x)+(3)(−1)=4x+5(Distribute) 6x+−3=4x+5 6x−3=4x+5 Step 2: Subtract 4x from both sides. 6x−3−4x=4x+5−4x 2x−3=5 Step 3: Add 3 to both sides. 2x−3+3=5+3 2x=8 Step 4: Divide both sides by 2. 2x 2 = 8 2 x=4 x=4 is the answer. also how many accounts do you have Cayden bruh..- \(\color{#0cbb34}{\text{End of Quote}}\) welp there u go

IMURDADDY4EVER:

Parentheses Exponents Multiplication and Division (from left to right) Addition and Subtraction (from left to right)

IMURDADDY4EVER:

So your going to do Parentheses first

supie:

So the first step would be to Distribute \(3(2x−1)=4x+5\) \((3)(2x)+(3)(−1)=4x+5\) \(6x+−3=4x+5\) \(6x−3=4x+5\) Then next we would have to get \(4x\) and subtract it from both of the sides of the equation \(6x−3−4x=4x+5−4x\) \(2x−3=5\) Then get the 3 and add it to both sides of the equation So \(+3\) to both sides \(2x−3+3=5+3\) \(2x=8\) Next we divide both sides by 2 (because it says \(2x\)) \(\LARGE\frac{2x}{2}=\frac{8}{2}\) So then \(\frac{2x}{2}\) cancels So all you have left is \(\frac{8}{2}\) Simplify that then you'll have your answer \(x=?\)

IMURDADDY4EVER:

well there you go. There's the answer

caydenisdead:

oh uh soo whats the answer

supie:

I told you what you have to do to get the answer You have to simplify \(\frac{8}{2}\)

sxdsouls:

Cayden work it out, I already gave it please review it correctly.

caydenmae:

ok